OpenStudy (anonymous):

25y^2-150y-4x^2-16x=-109 Put this hyperbola equation into standard form.

OpenStudy (anonymous):

Put x=0 and solve for y, you get the maximum and minimum points in the y axis, do the same thing for x, then, just put it in the formula:\[\left(\frac{ x-x_{0} }{ a }\right)^2+\left(\frac{ y-y _{0} }{ b }\right)^2=1\]where a=distance between the center and the extremum on the x axis, b is the same thing but for the y axis, and (x0, y0) is the center of the hyperbola. There probably is another way of doing it just by changing the equation, but I found it easier this way.

OpenStudy (anonymous):

thats sounds very complicated

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

Oh sorry, that is wrong, I was thinking about an ellipse

OpenStudy (anonymous):

oh ok thats ok